3) A wire loop of radius r = 2.00 mm has 80 turns, and 2.00 A of current is flowing through it. What is the maximum possible value of the torque that a magnetic field B = 2.50 T can exert on the loop?
Answer #1

We know that torque is given by:

Torque = N*i*A*B*sin

A = pi*R^2 = pi*(2.00*10^-3)^2 = 1.26*10^-5 m^2

R = Radius of loop = 2.00 mm = 2.00*10^-3 m

B = Magnetic field = 2.50 T

For maximum torque, sin = 1

N = number of loops = 80 turns

i = current in loop = 2.00 A

So,

Max torque = 80*2.00*1.26*10^-5*2.50*1

Max torque = 5.04*10^-3 N-m

Please Upvote.